| Operation Summary for December 26, 2011 - January 2, 2012 |
|
NuMI reported receiving 1.0E19 protons on target for the period from 12/26/11 to 1/2/12. Wednesday December 28 The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, and with the MiniBooNE and NuMI taking beam. During the day shift, Operations established beam to Meson FTBF T-1019 experiment at 10 AM. A MI abort-line vacuum pump (I:RV001) came into alarm around 10 AM. MI, Vacuum, and Controls personnel reported that the problem was in the tunnel, but that an access could wait until sometime next week. A Booster person conducted aperture scan studies and a Muon Rings person also conducted a study. During the evening shift, Operations disabled beam to T-1019 at 10:02 PM. Operators valved out a low volume LCW leak at the Tevatron sector A4 service building. Operators reported that a MI-39 kicker tripped off; it reset. Operators also reported that there was a minor vacuum burst in the P1 transfer line; everything reset. Thursday December 29 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, the NuMI decay pipe upstream LCW supply temperature came into alarm; Water Group personnel investigated. Operations reported that a Linac quadrupole power supply (L:QPS107) started tripping off when it hit its lower limit; Operators raised its current up by half an amp. Operators tuned MI and Linac. During the day shift, the Booster West Anode power supply tripped off; Booster personnel thought that a Booster RF station (BRF17) was to blame so they took it off line, repaired it, and brought it back on line within a half hour. Operations established beam to T-1019 at 10:45 AM. Muon Rings personnel began a study at 11 AM. During the evening shift, Muon Rings personnel completed their study. Operations disabled T-1019 beam at 10 PM. Operators tuned the I- Source, the 750 KeV Line, Booster, and MI and reported making some small improvements to the efficiencies. Friday December 30 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operators monitored the accelerator complex. During the day shift, at 8:26 AM, the MiniBooNE beam permit tripped off for the first of six times in 20 minutes due to a Switch magnet’s power supply current (MBEX) dropping low; Operators contacted a MiniBooNE person who advised lowering the trip limit. Operation established beam to T-1019 at 10:02 AM. During the evening shift, Operators reported that the MiniBooNE beam permit again tripped off on MBEX’s current drifting low; a MiniBooNE person told Operators to increase the MBEX voltage setting. Operators tuned the I- Source. Operations disabled T-1019 beam at 10 PM. Saturday December 31 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operations monitored the accelerator complex. During the day shift, Operators continued to increase the MBEX voltage setting. Operation established beam to T-1019 at 9:57 AM. Operators rebooted a NuMI BPM (NBPM65) at 12:39 PM. During the evening shift, Operations disabled T-1019 beam around 10 PM. After continuing to increase the MBEX voltage only to have the power supply to trip off anyway, Operations turned off beam to MiniBooNE and arranged for Power Supply personnel to come in on Sunday morning to investigate. Sunday January 1 The midnight shift began with Operations sending beam to NuMI. During the midnight shift, Operations monitored the accelerator complex. During the day shift, Power Supply personnel investigated the MBEX power supply, but found no problems, except that the voltage was set too high when they turned the supply back on. Operations resumed sending beam to MiniBooNE at 9:06 AM. Operation established beam to T-1019 at 10:48 AM. During the evening shift, Operations disabled T-1019 beam around 10 PM. Monday January 2 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operators tuned the I- Source, Linac, and the 750 KeV line. Operations Coordinator (OpsCo) The OpsCo asked Operations to continue monitoring the accelerator complex. |
|
More Information Comments and Suggestions |
| last modified 01/03/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |